A simple random sample from a population with a normal distribution of 106 body temperatures has x =98.90 degrees and s =0.61 degrees. Construct a 95​% confidence interval estimate of the standard deviation of body temperature of all healthy humans.

Answer by rothauserc(4718)

sample size(n) is 106 body temperatures
:
degrees of freedom = 106 - 1 = 105
:
the mean of the sample is 98.90
:
standard deviation of the sample is 0.61
:
We use the chi-square distribution to construct the 95% confidence interval(CI) for the population standard deviation
:
alpha(a) = 1 -(95/100) = 0.05
:
critical probability(p*) = 1 -(a/2) = 0.975
:
chi-square(0.975) = 78.536
:
chi-square(0.025) = 135.247
:
population variance(0.975) = ((106-1)*(0.61)^2)/78.536 = 0.4975
:
population variance(0.025) = ((106-1)*(0.61)^2)/135.247 = 0.2889
:
take the square roots to get standard deviation
:
************************************************
the 95% confidence interval is (0.5375, 0.7053)
************************************************
: